OpenStudy (anonymous):

While shopping at a clearance sale, Samantha finds a $60.00 dress on sale for 25% off. Samantha also has a 50% off coupon. Which statement correctly summarizes her savings? Which statement correctly summarizes her savings? Samantha will save $45.00 because the two sales will combine for 75% off the original cost. Samantha will save $37.50 because she must first find the 25% sale price before taking the extra 50% reduction. Samantha will save $22.50 because she must first find the 25% sale price before taking the extra 50% reduction.

OpenStudy (anonymous):

also she could save 15.00 but I don't think that's right

OpenStudy (dan815):

ok ill help you, so whats confusing you

OpenStudy (dan815):

okay so first you need to see how much the dress is 25% off and then 50% of that so 60*(1-0.25) *0.5 =? ^---50% off ^--25% off off

OpenStudy (dan815):

that is the new price of the dress

OpenStudy (dan815):

does it make sense?? 25% off means u are left with 75% of the price soo 60*0.75 = price after the 25% off 60*0.75 * 0.5 = price after the 50% off is applied to the 25% off price

OpenStudy (dan815):

60*0.75 * 0.5 = price after the 50% off is applied to the 25% off price 60*0.75 * 0.5 = 22.5 therefore she paid 22.5 for the new dress when it originally was 60 so she saved 60-22.5=37.5